OpenStudy (anonymous):

Can someone please help? tan(θ)+cot(θ)=(sec(θ))/(sin(θ)) L.S (sin^2(θ)+cos^2(θ))/(cos(θ))(sin(θ)) =1/(cos(θ))(sin(θ)) R.S sec(θ)/sin(θ) =1/cos(θ)/sin(θ) Please help. I always get stuck :(

OpenStudy (mertsj):

\[\tan \theta +\cot \theta =\frac{\sin \theta}{\cos \theta}+\frac{\cos \theta}{\sin \theta}=\frac{\sin ^2\theta+\cos ^2\theta}{\sin \theta \cos \theta}=\frac{1}{\sin \cos \theta}\]

OpenStudy (mertsj):

\[\frac{1}{\sin \theta \cos \theta}=\frac{1}{\sin \theta}\frac{1}{\cos \theta}=\frac{1}{\sin \theta}\sec \theta =\frac{\sec \theta}{\sin \theta}\]

OpenStudy (anonymous):

Does this mean that they are not an identity??

OpenStudy (mertsj):

What is the right side?

OpenStudy (anonymous):

the right side is sec(θ)/sin(θ)

OpenStudy (mertsj):

What did I show the left side is equal to?

OpenStudy (anonymous):

1/sin(θ)cos(θ)

OpenStudy (mertsj):

Look at the window right below that where the work continues.

OpenStudy (anonymous):

oh okay. Thank you very much

OpenStudy (mertsj):

yw